Grade 12th passTrigonometry

3sinx+4cosx=5 answer the question and my friend to Z

Ashish Rana
9 Years agoGrade 12th pass

1 Answer

Nishant Vora
9 Years ago
3/5sinx+4/5cosx=1

let cosA = 3/5 hence sinA=4/5

so now
cosAsinx+sinAcosx=1
sin(A+x)= sin(pi/2)

A + x = npi + (-1)^n pi/2